# Quantum geometric formulation of Brans-Dicke theory for Bianchi I spacetime

### v2

@@ III Symmetry Reduction of BDT for the Bianchi I model
The Bianchi Class A model, in general, represents all possible homogeneous spacetimes with different degrees of symmetry. The four-dimensional Bianchi spacetime has a topology of the $\Sigma\times\mathbb{R}$, where the intrinsic metric $q_{ab}$ defined on the spatial manifold $\Sigma$ is homogeneous with three-dimensional symmetry group that is simply transitive. The symmetry of the spatial hypersurface is determined by the Killing vectors. Essentially, the killing vector fields give the direction along which Lie dragging the metric does not change the metric [45]. The fiducial triads and co-triads, denoted, respectively, by $\mathring{e}^{a}_{i}$ and $\mathring{e}^{i}_{a}$, satisfy the relations $\mathring{e}^{a}_{i}\mathring{e}^{i}_{b}=\delta^{a}_{b}$ and $\mathring{e}^{a}_{i}\mathring{e}^{j}_{a}=\delta^{j}_{i}$. The geometry of Bianchi I, which is one of the simplest types of Bianchi Class A model, is such that it admits the universe to have different scale factors in each principal direction, but with zero spatial curvature.
@@
It is to be noted that a fiducial triad is introduced here at each point of the spatial manifold to carry out the integration as $\Sigma$ is non-compact and they commute with the killing vector fields denoted by $\mathring{\xi}^{a}_{i}$, $[\mathring{e}_{i},\mathring{\xi}_{j}]^{a}=0$. The underlying structure of the spatial hypersurface is dictated by the commutator relations, given by $[\mathring{e}_{i},\mathring{e}_{j}]^{a}=C^{k}_{ij}\mathring{e}^{a}_{k}$. The values of the structure constants $C^{k}_{ij}$ completely define all the symmetry of the spatial sector. The Bianchi I model considered in this work is specified by $C^{i}_{jk}=0$. To be specific, the Bianchi I hypersurface $\Sigma$ is flat and, therefore, $\Gamma^{i}_{a}=0$. Also, the homogeneity implies the contribution to the Hamiltonian from the spatial derivative is identically zero. This, in turn, it greatly simplifies the expression of the curvature tensor. The diagonal Bianchi I universe in terms of the Asthekar variables and densetized triads are parameterized by
$A^{i}_{a}$ $=$ $\tilde{c_{i}}V_{o}^{-1/3}\mathring{e}^{a}_{i},$ (17)
$E^{b}_{j}$ $=$ $p_{j}V_{o}^{-2/3}\sqrt{q}\mathring{e}^{b}_{j},$ (18)
@@
where there is no summation over indexes $i$ or $j$. Let us note that Bianchi I is a homogeneous but anisotropic spacetime. Therefore, $\tilde{c_{i}}$ and $p_{j}$ only depend on time. Also, with the information that there is no spatial curvature for Bianchi I, $\Gamma^{i}_{a}=0$, we shall show that the only constraint that survives off-shell is the Hamiltonian constraint. Below, we investigate the Gauss, diffeomorphism and Hamiltonian constraints in the given context.
The Gauss constraint is
$\mathcal{G}_{i}=D_{a}E^{a}_{i},$ (19)
which is identically zero because Bianchi-I is a homogeneous and, therefore, the derivative with respect to spatial coordinate vanishes identically. The diffeomorphism constraint is
$\mathcal{V}_{a}$ $=$ $\frac{1}{\gamma}F^{i}_{ab}E^{b}_{i}=\frac{1}{\gamma}\epsilon^{i}_{kl}A^{k}_{a}A^{l}_{b}E^{b}_{i}\propto\epsilon^{i}_{kl}\delta^{l}_{i}.$ (20)
Therefore, the diffeomorphism constraints vanish identically for the Bianchi I model. Thus, we are only left with the scalar (or Hamiltonian) constraint, which is given by
@@
$+$ $\frac{1}{3+2\omega}\left[\frac{(\tilde{K}^{i}_{a}E^{a}_{i})^{2}}{\phi\sqrt{q}}+2\frac{(\tilde{K}^{i}_{a}E^{a}_{i})\pi_{\phi}}{\sqrt{q}}+\frac{\pi^{2}_{\phi}}{\sqrt{q}}\right]$
@@
Now, substituting the curvature tensor, $F^{j}_{ab}=\epsilon^{i}_{kl}(\Gamma^{k}_{a}+\gamma\tilde{K}^{k}_{a})(\Gamma^{l}_{b}+\gamma\tilde{K}^{l}_{b})$, due to the fact that Bianchi I is spatially flat, we find that $\Gamma^{k}_{a}=0$ and, hence, we are left with
@@
We also notice that
@@
where $[a,b]\equiv(ab-ba)/2$ denotes the antisymmetrization, with which, the expression for $\mathcal{H}$ can be further reduced to
$\mathcal{H}$ $=$ $-\frac{1}{\phi\sqrt{q}}\left[E^{a}_{m}E^{b}_{n}\tilde{K}^{m}_{[a}\tilde{K}^{n}_{b]}+\right.$ (24)
-- $\left.\frac{1}{(3+2\omega)}\left(\tilde{K}^{i}_{a}E^{a}_{i}+\,\pi_{\phi}\phi\right)^{2}\right].$
In the limiting case of $\omega\to\infty$ and $\phi$ being a constant, the scalar constraint of the BDT gets reduced to the scalar constraint of Einstein’s GR in the vacuum.
The next important step is to cast the theory in terms of the Asthekar’s variables. The canonical structure is given by
$\{A^{i}_{a}(x),E^{b}_{j}(y)\}=\gamma\delta^{i}_{j}\delta^{b}_{a}\delta(x-y).$ (25)
It is to be noted that the Bianchi-I spacetime is non-compact and in order to carry out the integrations, we consider a fiducial volume $V_{0}$. Evaluating each term of the Hamiltonian individually by using the parameterization of Eq. (17), we have
$E^{a}_{m}E^{b}_{n}\tilde{K}^{m}_{[a}\tilde{K}^{n}_{b]}$ $=$ $\frac{\mathring{q}}{\gamma^{2}l_{0}^{6}}\left(\tilde{c}_{1}\tilde{c}_{2}p_{1}p_{2}+\tilde{c}_{2}p_{2}\tilde{c}_{3}p_{3}\right.$ (26)
$+$ $\left.\tilde{c}_{1}p_{1}\tilde{c}_{3}p_{3}\right),$
$\tilde{K}^{i}_{a}E^{a}_{i}$ $=$ $\frac{\sqrt{\mathring{q}}}{\gamma l_{0}^{3}}(\tilde{c}_{1}p_{1}+\tilde{c}_{2}p_{2}+\tilde{c}_{3}p_{3}).$ (27)
Now, combining all the terms and using that $\sqrt{q}=\sqrt{\mathring{q}}\sqrt{p_{1}p_{2}p_{3}}/l_{0}^{3}$ in Eq. (24), we find
@@
where we have redefined the conjugate momentum corresponding to the scalar field as $\pi_{\phi}\equiv p_{\phi}\sqrt{\mathring{q}}/l_{0}^{3}$ and $\beta=3+2\omega$. The smeared Hamiltonian is obtained by integrating the symmetry reduced Hamiltonian, Eq. (28), over the fiducial volume $V_{0}$, leading to
@@
where we have used that $\int d^{3}x\sqrt{\mathring{q}}=l_{0}^{3}$. The Poisson flow of a Dirac observable $\dot{\mathcal{O}}=\{\mathcal{O},C_{H}\}$ with the Hamiltonian gives the equation of motion.
IV Classical dynamics of BDT in Bianchi-I model
@@
The dynamics is given by the Poisson flow of the variable with the smeared Hamiltonian. Then, the equations of motion for $c_{I}$ and its conjugate variables $p_{I}$ are
$\dot{\tilde{c}}_{I}$ $=$ $\gamma\,\delta_{IJ}\frac{\delta C_{\mathcal{H}}}{\delta p_{J}},$ (32)
$\dot{p_{I}}$ $=$ $-\gamma\,\delta_{IJ}\frac{\delta C_{\mathcal{H}}}{\delta\tilde{c}_{J}},$ (33)
@@
$\dot{\phi}$ $=$ $\frac{\delta C_{\mathcal{H}}}{\delta\pi_{\phi}},$ (34)
$\dot{\pi}_{\phi}$ $=$ $-\frac{\delta C_{\mathcal{H}}}{\delta\phi}.$ (35)
The cosmic time interval can be written as $dt=Ndt^{\prime}$, whereas the time variable $t^{\prime}$ will be determined by the choice of the lapse function $N$. In this work, we choose $N=\sqrt{p_{1}p_{2}p_{3}}$, then we have
$dt=\sqrt{p_{1}p_{2}p_{3}}dt^{\prime}.$ (36)
Using the Hamiltonian constraint, Eq. (29), we can now derive the equations of motion. These will be obtained with respect to the time $t^{\prime}$ using Eq. (36). From Eq. (32), we obtain the following equations of motion for the $c_{I}$ variables,
$\tilde{c}_{I}^{\prime}$ $=$ $-\frac{\tilde{c}_{I}}{\gamma\phi}\Bigl{[}\tilde{c}_{J}p_{J}+\tilde{c}_{K}p_{K}$
@@
where refers to the derivative with respect to $t^{\prime}$. Similarly, from Eq. (33), we obtain the equations of motion for the conjugate variables $p_{I}$,
${p}_{I}^{\prime}$ $=$ $\frac{p_{I}}{\gamma\phi}\Bigl{[}\tilde{c_{J}}p_{J}+\tilde{c}_{K}p_{K}$
@@
The set of Eqs. (IV)-(40) completely specifies the dynamics of BDT in the Bianchi-I spacetime at the classical level. Also, even at the level of equations of motion, it is straightforward to see that in the limit $\omega\rightarrow\infty$ and constant $\phi$, the equations of motion for the Einstein-Hilbert action are recovered [45].
V Quantum Corrected Dynamics of BDT in Bianchi-I Spacetime
@@
The quantum corrected effective dynamics of LQC have been studied in detail in [46], which incorporate the leading-order quantum geometric effects [47]. These effective models work very well in comparison with the full quantum dynamics of LQC even in the deep quantum regime [20], especially for the states that are sharply peaked on a classical trajectory at late times [48]. In the homogeneous and isotropic FLRW universe, it is obtained by the replacement $c\rightarrow\sin\left(\mu c\right)/\mu$, where $c$ denotes the conjugate momentum of the area operator $p$ ($\propto a^{2}$, where $a$ is the expansion factor of the universe), and the lattice spacing, $\mu$, is called the polymerization parameter. Different choices of the parameter $\mu$ give rise to different schemes of quantization with distinct effective dynamics. The expectation is that this would act as an ultraviolet regulator and mitigate the initial singularity [49, 50].
In the literature, various choices of $\mu$ have been considered [20, 51]. In the original scheme, the area gap was implemented as a kinematic feature on the comoving frame, which is equivalent to introducing a fixed lattice of constant spacing. On the other hand, the improved scheme is equivalent to introducing the minimum area eigenvalue on the physical frame, in contrast to the original. In this paper we shall adopt two different $\bar{\mu}$ schemes in order to explore different aspects of the model. We denominate them as $\bar{\mu}$ type A and B, to be denoted as $\bar{\mu}_{A}$ and $\bar{\mu}_{B}$, respectively, for our purpose. Next, we describe these two quantization schemes in more details.
V.1 The $\bar{\mu}_{A}$ quantization scheme
@@
$\bar{\mu}_{I}$ $=$ $\sqrt{\frac{\Delta}{p_{I}}},$ (41)
where $\Delta=4\pi\sqrt{3}\gamma l_{Pl}^{2}$ is the area gap in the full theory of LQC.
Clearly, when $p_{I}\gg\Delta$, we have $\bar{\mu}_{I}\rightarrow 0$, and expect that the quantum effects are very small, and the classical limit is obtained. However, near the singular point, $p_{I}\simeq 0$, we have $\bar{\mu}_{I}\gg 1$, such that the quantum effects are expected to be very large. In the following, we shall consider whether such effects are larger enough such that the spacetime singularity that used to appear at $p_{I}=0$ will now be regulated in the current setup.
@@
In order to study the effective theory of loop quantum Brans-Dicke cosmology, we also want to know the effect of matter fields on the dynamical evolution. Hence, we have included in Eq. (42) an extra massless scalar matter field $\Phi$, with constant conjugate momenta $P_{\Phi}$, which acts as a clock and has energy density $\rho=P_{\Phi}^{2}/(2p_{1}p_{2}p_{3})$. Quantum corrected equations of motion for the symmetry reduced connection parameterized by $\tilde{c}$ for the $\bar{\mu}_{A}$ scheme given by Eq. (41) are given by
@@
-- $\left.\frac{2}{\beta}\left[\frac{\sin(\bar{\mu}_{I}\tilde{c}_{I})}{\bar{\mu}_{I}}p_{I}+\frac{\sin(\bar{\mu}_{J}\tilde{c}_{J})}{\bar{\mu}_{J}}p_{J}+\frac{\sin(\bar{\mu}_{K}\tilde{c}_{K})}{\bar{\mu}_{K}}p_{K}+\gamma\pi_{\phi}\phi\right]\right\}$
@@
For other choices of schemes for $\bar{\mu}_{I}$, the change will be in the terms between the first square brackets. Similarly, quantum corrected equations of motion for the symmetry reduced densetized triads parameterized by $p_{I}$ are given by
@@
-- $\left.\frac{2}{\beta}\left[\frac{\sin(\bar{\mu}_{I}\tilde{c}_{I})}{\bar{\mu}_{I}}p_{I}+\frac{\sin(\bar{\mu}_{J}\tilde{c}_{J})}{\bar{\mu}_{J}}p_{J}\right.\right.$
@@
The set of Eqs. (43)-(V.1) completely define the quantum corrected dynamics of the gravitational sector of the Bianchi-I spacetime for BDT in the Jordan’s frame in $\bar{\mu}_{A}$ scheme. At this point, we can compute the average Hubble parameter $H$ for the quantum Bianchi-I spacetime,
$H\equiv\frac{H_{1}+H_{2}+H_{3}}{3},$ (47)
where we have defined the average scale factor as $a=\sqrt{a_{1}a_{2}a_{3}}$, $p_{I}=a_{J}a_{K}l_{0}^{2}$ for non repeated indexes $I,\,J,\,K$ and $H_{I}=\dot{a}_{I}/a_{I}$. The average Hubble parameter reads
$H$ $=$ $\frac{1}{6\gamma\phi\sqrt{p_{1}p_{2}p_{3}}}\left\{\left[\frac{\sin(\bar{\mu}_{2}\tilde{c}_{2})}{\bar{\mu}_{2}}p_{2}+\frac{\sin(\bar{\mu}_{3}\tilde{c}_{3})}{\bar{\mu}_{3}}p_{3}\right]\cos(\bar{\mu}_{1}\tilde{c}_{1})\right.$
@@
-- $\frac{1}{6\sqrt{p_{1}p_{2}p_{3}}}\frac{\dot{\phi}}{\phi}\left[\cos(\bar{\mu}_{1}\tilde{c}_{1})+\cos(\bar{\mu}_{2}\tilde{c}_{2})+\cos(\bar{\mu}_{3}\tilde{c}_{3})\right],$
where $\sqrt{p_{1}p_{2}p_{3}}$ comes from Eq. (36). It should be noted that at each point of the evolution of the universe dictated by Eq. (LABEL:HubbleParameter), the condition for non-singular quantum bounce, $H=0$, holds when we impose that $p_{1},p_{2},p_{3}\neq 0$ simultaneously. One can notice from Eq. (LABEL:HubbleParameter) that this condition will be satisfied only for $\cos(\bar{\mu}_{I}\tilde{c}_{I})=0$ for $I=1,2,3$. However, this is a stringent condition as $\tilde{c}_{I}$ are all independent and one can not guarantee all of them to be zero simultaneously. In order to guarantee that a non-singular behavior happens indeed, we now follow the perspective of Ref. [36].
@@
Now, one remembers that $\rho=P_{\Phi}^{2}/(2p_{1}p_{2}p_{3})$ is the energy density for a massless scalar field, then $P_{\Phi}$ is a constant of motion. Therefore, the term $\left(\rho+p_{I}\frac{\partial\rho}{\partial p_{I}}\right)$ in Eq. (50) vanishes, which means that $\mathcal{G}_{I}$ are constants. For constant $\mathcal{G}_{I}$, from Eq. (49), given that the sines have a maximum value of unit, it sets a lower bound on $p_{I}$,
@@
V.2 The $\bar{\mu}_{B}$ quantization scheme
In order to express the average Hubble parameter in a closed form and to see the singularity resolution more directly, the use of the alternative $\bar{\mu}_{B}$ quantization scheme becomes more appropriate. In the $\bar{\mu}_{B}$ quantization scheme, we have that [40]

### v3

@@ III Symmetry Reduction of Classical BDT for Bianchi I Spacetimes
have a topology of $\Sigma\times\mathbb{R}$, where $N$ and $a_{I}$ are functions of $t$ only, $x^{\mu}\equiv(t,x^{a})$ and $x^{a}\equiv(x,y,z)$. Thus, the intrinsic metric $q_{ab}=\text{diag}(a_{1}^{2},a_{2}^{2},a_{3}^{2})$ defined on the spatial manifold $\Sigma$ is homogeneous with three-dimensional symmetry group that is simply transitive. The symmetry of the spatial hypersurface is determined by the Killing vectors. Essentially, the Killing vector fields give the direction along which the metric does not change under the Lie dragging [43].
III.1 Classical Hamiltonian
Considering a fiducial cell with lengths $L_{1},\;L_{2}$ and $L_{3}$ along the three spatial directions [32],
$ds_{o}^{2}\equiv{}^{o}q_{ab}dx^{a}dx^{b}=dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2},$ (15)
its triads and cotriads, denoted, respectively, by $\mathring{e}^{a}_{i}$ and $\mathring{e}^{i}_{a}$, satisfy the relations $\mathring{e}^{a}_{i}\mathring{e}^{i}_{b}=\delta^{a}_{b}$ and $\mathring{e}^{a}_{i}\mathring{e}^{j}_{a}=\delta^{j}_{i}$. Without loss of the generality, one often sets $L_{1}=L_{2}=L_{3}\equiv l_{0}$, which will be adopted in this paper. The geometry of the Bianchi I spacetimes is one of the simplest types of Bianchi Class A models [63], such that it admits the Universe to have different scale factors in each principal direction, but with zero spatial curvature.
@@ III.1 Classical Hamiltonian
The introduction of fiducial triads trivializes the gauge freedom corresponding to the Gauss and the spatial diffeomorphisms constraints and they commute with the Killing vector fields denoted by $\mathring{\xi}^{a}_{i}$, satisfying the relations $[\mathring{e}_{i},\mathring{\xi}_{j}]^{a}=0$. The underlying structure of the spatial hypersurface is dictated by the commutator relations, given by $[\mathring{e}_{i},\mathring{e}_{j}]^{a}=C^{k}_{ij}\mathring{e}^{a}_{k}$. The values of the structure constants $C^{k}_{ij}$ completely define all the symmetry of the spatial sector. The Bianchi I model considered in this work is specified by $C^{i}_{jk}=0$. To be specific, the Bianchi I hypersurface $\Sigma$ is flat. Also, the homogeneity implies the contribution to the Hamiltonian from the spatial derivative is identically zero. The diagonal Bianchi I Universe in terms of the Asthekar variables and densetized triads are parametrized by
$A^{i}_{a}=\tilde{c_{i}}V_{o}^{-1/3}\mathring{e}^{a}_{i},\quad E^{b}_{j}=p_{j}V_{o}^{-2/3}\sqrt{q}\mathring{e}^{b}_{j},$ (16)
which satisfy the canonical relations
$\{A^{i}_{a}(x),E^{b}_{j}(y)\}=\gamma\delta^{i}_{j}\delta^{b}_{a}\delta(x-y),$ (17)
@@
where there is no summation over indexes $i$ or $j$, and the volume of the fiducial cell $V_{o}$ is given by $V_{o}=l_{0}^{3}$. Since Bianchi I spacetimes are homogeneous but anisotropic, $\tilde{c_{i}}$ and $p_{j}$ depend only on time. Also, since there is no spatial curvature, we have $\Gamma^{i}_{a}=0$. As a result, the only constraint that survives after the adoption of the fiducial triad is the Hamiltonian constraint, which is given by
@@
$+$ $\frac{1}{3+2\omega}\left[\frac{(\tilde{K}^{i}_{a}E^{a}_{i})^{2}}{\phi\sqrt{q}}+2\frac{(\tilde{K}^{i}_{a}E^{a}_{i})p_{\phi}}{\sqrt{q}}+\frac{p^{2}_{\phi}\phi}{\sqrt{q}}\right]$
@@
Since the Bianchi I spacetimes are spatially flat, $\Gamma^{k}_{a}=0$, as pointed out, we find that
@@
We also notice that for the parametrization  (16), we have
@@
$E^{a}_{m}E^{b}_{n}\tilde{K}^{m}_{[a}\tilde{K}^{n}_{b]}=\frac{\mathring{q}}{\gamma^{2}l_{0}^{6}}\left(\tilde{c}_{1}\tilde{c}_{2}p_{1}p_{2}+\tilde{c}_{2}p_{2}\tilde{c}_{3}p_{3}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\left.\tilde{c}_{1}p_{1}\tilde{c}_{3}p_{3}\right),$ (21)
$\tilde{K}^{i}_{a}E^{a}_{i}=\frac{\sqrt{\mathring{q}}}{\gamma l_{0}^{3}}(\tilde{c}_{1}p_{1}+\tilde{c}_{2}p_{2}+\tilde{c}_{3}p_{3}),\quad$ (22)
where $[a,b]\equiv(ab-ba)/2$ denotes the antisymmetrization. In the limiting case of $\omega\to\infty$ and $\phi$ being a constant, the scalar constraint of the BDT gets reduced to the scalar constraint of Einstein’s GR in the vacuum.
Now, combining all the terms and using that $\sqrt{q}=\sqrt{\mathring{q}}\sqrt{p_{1}p_{2}p_{3}}/l_{0}^{3}$, we finally obtain
@@
where, in the move to homogeneous variables, we have redefined the conjugate momentum corresponding to the scalar field as $p_{\phi}\equiv\pi_{\phi}\sqrt{\mathring{q}}/l_{0}^{3}$ and $\beta=3+2\omega$. The smeared Hamiltonian is obtained by integrating the symmetry reduced Hamiltonian, Eq. (23), over the fiducial volume $V_{0}$, leading to
@@
where we have used $\int d^{3}x\sqrt{\mathring{q}}=l_{0}^{3}$.
III.2 Classical dynamics
@@
The dynamics of a system is given by the Poisson equation, $\dot{\mathcal{O}}=\{\mathcal{O},C_{H}\}$ for the phase space function $\mathcal{O}$. Then, the equations of motion for $c_{I}$ and its conjugate variables $p_{I}$ are
${\tilde{c}}^{\prime}_{I}$ $=$ $\gamma\,\delta_{IJ}\frac{\delta C_{\mathcal{H}}}{\delta p_{J}},$ (27)
${p_{I}}^{\prime}$ $=$ $-\gamma\,\delta_{IJ}\frac{\delta C_{\mathcal{H}}}{\delta\tilde{c}_{J}},$ (28)
@@
${\phi}^{\prime}$ $=$ $\frac{\delta C_{\mathcal{H}}}{\delta\pi_{\phi}},$ (29)
${\pi}_{\phi}^{\prime}$ $=$ $-\frac{\delta C_{\mathcal{H}}}{\delta\phi},$ (30)
where a dot denotes the derivative with respect to cosmic time $t$ and a prime denotes derivative with respect to $\tau$ defined as
$d\tau\equiv Ndt=\sqrt{p_{1}p_{2}p_{3}}dt.$ (31)
In this work, we choose $N=\sqrt{p_{1}p_{2}p_{3}}$.
Using the Hamiltonian constraint, Eq. (24), we can now derive the equations of motion. In particular, from Eq. (27), we obtain the following equation of motion for the $c_{I}$ variables,
$\tilde{c}_{I}^{\prime}$ $=$ $-\frac{\tilde{c}_{I}}{\gamma\phi}\Bigl{[}\tilde{c}_{J}p_{J}+\tilde{c}_{K}p_{K}$ (32)
@@
where $J,\;K\not=I$.
Similarly, from Eq. (28), we obtain the equation of motion for the conjugate variables $p_{I}$,
${p}_{I}^{\prime}$ $=$ $\frac{p_{I}}{\gamma\phi}\Bigl{[}\tilde{c_{J}}p_{J}+\tilde{c}_{K}p_{K}$ (33)
@@
The set of Eqs. (32)-(35) completely specifies the classical dynamics of the Bianchi-I spacetimes for BDT. Note that even at the level of equations of motion, it is straightforward to see that in the limit $\omega\rightarrow\infty$ and constant $\phi$, the equations of motion for the Einstein-Hilbert action are recovered [43].
IV Loop Quantum Effective Dynamics of Bianchi I Spacetimes in BDT
The loop effective quantum dynamics for homogeneous and isotropic Universe in LQC [64] have been studied extensively, and now it is well established [12] that the effective dynamics incorporate the leading-order quantum geometric effects of the full theory very well even in the deep Planck regime [65]. This is especially true for the states that are sharply peaked on a classical trajectory at late times [66].
In the homogeneous and isotropic FLRW Universe, it is obtained by the replacement
$c\rightarrow\frac{\sin\left(\mu c\right)}{\mu},$
@@ IV Loop Quantum Effective Dynamics of Bianchi I Spacetimes in BDT
where $c$ denotes the conjugate momentum of the area operator $p$ ($\propto a^{2}$, where $a$ is the expansion factor of the Universe), and the lattice spacing, $\mu$, is called the polymerization parameter. Different choices of the parameter $\mu$ give rise to different schemes of quantization with distinct effective dynamics. The expectation is that this would act as an ultraviolet regulator and mitigate the initial singularity [67, 68].
In the literature, various choices of $\mu$ have been considered [12, 69] 111 It should be noted that the choices of $\mu$ are purely geometric and are independent of the underlying theory. As a result, they are applicable to the Brans-Dicke theory considered in this paper. For more details, see Ref. [70] and references therein.. In particular, in the original scheme–the $\mu_{o}$ scheme [71], the area gap was implemented as a kinematic feature on the comoving frame, which is equivalent to introducing a fixed lattice of constant spacing, so that the area $\Box_{ij}$ in the ($i,j$) plane is measured by the fiducial metric ${}^{o}q_{ab}$ defined by Eq. (15). However, it was found that this leads to inconsistencies [72]. After extensive investigations in the past two decades or so (see, for example, Ref. [69] and references therein), it was found that the $\bar{\mu}$ scheme [73], which is sometimes also referred to as improved dynamics, is the unique one that leads to consistent physics not only for the homogeneous and isotropic quantum background, but also for its linear perturbations [12, 13, 15, 16]. The distinguishable feature of the $\bar{\mu}$ scheme comes from the observation that the quantization of area should refer to physical geometries instead of the fiducial metric. In particular, the area $\Box_{ij}$ in the ($i,j$) plane should be measured by the physical metric $q_{ab}$ defined in Eq. (14), instead of the fiducial metric ${}^{o}q_{ab}$ defined by Eq. (15). Since the physical area of the faces of an elementary cell is $p$, and each side of $\Box_{ij}$ is $\sqrt{\Delta}$ times the edge of the elementary cell, we are led to [73]
$\bar{\mu}^{2}p=\Delta\quad\Rightarrow\quad\bar{\mu}=\sqrt{\frac{\Delta}{p}},$ (36)
where $\Delta\equiv 4\pi\sqrt{3}\gamma l_{Pl}^{2}$ is the area gap in the full theory of LQG [7].
When generalizing the above $\bar{\mu}$ scheme to the homogeneous but anisotropic cases, including the Bianchi I models, in the literature there exist two different schemes [28, 29, 30, 32], which we shall introduce below and refer them to as the $\bar{\mu}_{A}$ and $\bar{\mu}_{B}$ schemes, respectively. Both schemes have been studied in great details [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], and showed that, among other interesting features, the singularities are resolved and replaced by quantum bounces. More recently  [47], it was found that the $\bar{\mu}_{A}$ quantization suffers from a problem in which one of the triad components and associated polymerized term retains Planckian character even at large volumes. As a result, not only is the anisotropic shear not preserved across the bounce, but also the Universe can exhibit an unexpected cyclic evolution.
IV.1 The $\bar{\mu}_{A}$ quantization scheme
@@
$\bar{\mu}_{I}$ $=$ $\sqrt{\frac{\Delta}{p_{I}}}.$ (37)
Clearly, when $p_{I}\gg\Delta$, we have $\bar{\mu}_{I}\rightarrow 0$, and expect that the quantum effects are very small and the classical limit is obtained. However, near the singular point, $p_{I}\simeq 0$, we have $\bar{\mu}_{I}\gg 1$, such that the quantum effects are expected to be very large. In the following, we shall consider whether such effects are larger enough so that the spacetime singularity that used to appear classically at $p_{I}=0$ will now be regulated in the current setup.
@@
Note that, in order to study the effective theory of loop quantum Brans-Dicke cosmology, we also want to know the effect of matter fields on the dynamical evolution. Hence, we have included in Eq. (38) an extra massless scalar matter field $\Phi$, with momentum conjugate $P_{\Phi}$, and the corresponding energy density is given by
$\rho=\frac{P_{\Phi}^{2}}{2p_{1}p_{2}p_{3}}.$ (39)
Since $C_{\rm eff}$ does not depend explicitly on the scalar field $\Phi$, we have
$P_{\Phi}^{\prime}=-\frac{\partial C_{\rm eff}}{\partial\Phi}=0,$ (40)
that is, $P_{\Phi}$ is a constant of motion. On the other hand, the other quantum corrected Hamilton’s equations for the symmetry reduced connection parametrized by $\tilde{c}$ with the $\bar{\mu}_{A}$ scheme given by Eq. (37) are given by
@@
$-\left.\frac{2}{\beta}\left[\frac{\sin(\bar{\mu}_{I}\tilde{c}_{I})}{\bar{\mu}_{I}}p_{I}+\frac{\sin(\bar{\mu}_{J}\tilde{c}_{J})}{\bar{\mu}_{J}}p_{J}+\frac{\sin(\bar{\mu}_{K}\tilde{c}_{K})}{\bar{\mu}_{K}}p_{K}+\gamma\pi_{\phi}\phi\right]\right\}$
@@
For other choices of schemes for $\bar{\mu}_{I}$, the change will be in the terms inside the first square brackets. Similarly, quantum corrected equations of motion for the symmetry reduced densetized triads parametrized by $p_{I}$ are given by
@@
$-\left.\frac{2}{\beta}\left[\frac{\sin(\bar{\mu}_{I}\tilde{c}_{I})}{\bar{\mu}_{I}}p_{I}+\frac{\sin(\bar{\mu}_{J}\tilde{c}_{J})}{\bar{\mu}_{J}}p_{J}\right.\right.$
@@
The set of Eqs. (41)-(IV.1) completely define the quantum corrected dynamics of the Bianchi-I spacetime in $\bar{\mu}_{A}$ scheme for BDT in the Jordan’s frame.
Once the above equations are solved, we can compute the average Hubble parameter $H$ and the expansion factor $a$ for the effective quantum Bianchi-I spacetime
$H\equiv\frac{H_{1}+H_{2}+H_{3}}{3},\quad a\equiv\sqrt{a_{1}a_{2}a_{3}},$ (45)
where $H_{I}=\dot{a}_{I}/a_{I}$ and $p_{I}=a_{J}a_{K}l_{0}^{2}$ for nonrepeated indexes $I,\,J,\,K$. In particular, the average Hubble parameter reads
$H$ $=$ $\frac{1}{6\gamma\phi\sqrt{p_{1}p_{2}p_{3}}}\left\{\left[\frac{\sin(\bar{\mu}_{2}\tilde{c}_{2})}{\bar{\mu}_{2}}p_{2}+\frac{\sin(\bar{\mu}_{3}\tilde{c}_{3})}{\bar{\mu}_{3}}p_{3}\right]\right.$
×$\times\cos(\bar{\mu}_{1}\tilde{c}_{1})$
@@
$-\frac{\dot{\phi}}{6\phi\sqrt{p_{1}p_{2}p_{3}}}\left[\cos(\bar{\mu}_{1}\tilde{c}_{1})+\cos(\bar{\mu}_{2}\tilde{c}_{2})+\cos(\bar{\mu}_{3}\tilde{c}_{3})\right],$
where the factor $\sqrt{p_{1}p_{2}p_{3}}$ comes from Eq. (31). To see how classical singularities can be avoided in the framework of LQG, let us provide some heuristic arguments, by following the analysis provided in  [29]. For more details, see our numerical analyses to be presented in the next section.
@@
Since $\rho$ is given by Eq. (39), we find the identity
$\rho+p_{I}\frac{\partial\rho}{\partial p_{I}}=0,$ (51)
which means that $\mathcal{G}_{I}$ are constants. For constant $\mathcal{G}_{I}$, from Eq. (47), given that the sines have a maximum value of unit, it sets a lower bound on $p_{I}$,
@@
IV.2 The $\bar{\mu}_{B}$ quantization scheme
In the $\bar{\mu}_{B}$ scheme, we have  [40]